OpenStudy (anonymous):

A man divided $9,000 among his wife, son, and daughter. The wife received twice as much as the daughter, and the son received $1,000 more than the daughter. How much did each receive? If x is the amount the wife received, then which of the following expressions represents the amount received by the son? + 1,000 x + 1,000 2x + 1,000

OpenStudy (anonymous):

if i deciphered the question right i would say x+1000

OpenStudy (anonymous):

well you think you can explain it? your prob right but i have some of the same type of questions and i dont want to come back here

OpenStudy (kkutie7):

the first answer for the son is + 1,000?

OpenStudy (kkutie7):

just from looking at the question I would tend to write things like this: x=wife, s=son, d= daughter. x=2*d, s=1,000+d. x/2=d. s=1,000+x/2

OpenStudy (anonymous):

alright thanks for the explanation :D and answer